OpenStudy (ayonnaleflore):

What is the second term in the expansion of (2x + 7)6? 1344x4 23,520x5 1344x5 23,520x4

OpenStudy (welshfella):

(r + 1)th term of (a + b)^n= nCr a^(n-r)b^r

OpenStudy (welshfella):

in this case r = 1, n = 6 , a = 2x and b = 7

OpenStudy (ayonnaleflore):

so (2x+7)^6= 6^(6-1)???

OpenStudy (welshfella):

no 6C2 (2x)^5 * 7^1

OpenStudy (welshfella):

Sorry that' s nC1 which = 6

OpenStudy (welshfella):

(2x)^5 = 32 x^5 and 7^1 = 7

OpenStudy (ayonnaleflore):

so 32x^5*7

OpenStudy (welshfella):

6C1 = 6

OpenStudy (welshfella):

you forgot the 6

OpenStudy (welshfella):

6C1 * 32x^5 * 7 = 6 * 32x^5 * 7

OpenStudy (ayonnaleflore):

32x^5 * 42

OpenStudy (welshfella):

right Now sismplify some more

OpenStudy (ayonnaleflore):

ok so how would i simlify it more ??

OpenStudy (welshfella):

just multiply 32 by 42

OpenStudy (ayonnaleflore):

1344x^5 right

OpenStudy (welshfella):

yes Thats the 3rd choice

OpenStudy (ayonnaleflore):

ok thank yo

OpenStudy (welshfella):

yw